A proton moves through a magnetic field of magnitude 4.5 T at a speed of 7.08×10^6 m/s


Perpendicular to the field. Find the


a. Centripetal acceleration



b. Radius of the circular path of the proton

1
Expert's answer
2022-05-24T13:56:50-0400

"qvB=ma,\implies a=\frac{qvB}m=5.6\cdot 10^{18}~\frac{m}{s^2},"

"r=\frac{mv}{qB}=9\cdot 10^{-6}~m."